Systems Staff Meeting Notes 07/28/03

Systems Staff Meeting Notes
July 28, 2003

Present: Cynthia Bowker, Will Chaney, Bill Covey, Suzy Covey, Debra Fetzer, Winston Harris, Bobbie Meng, Hilton Soto, Geoffrey Stewart, Ying Tang
  1. Copier – interrupt does not stop the current job, just lets you make a different copy then resumes the original job.
  2. Bad rain on Friday. Ying wore shopping bags on her feet to walk to her car.
  3. Rochester has some very foolish drivers.
  4. Library Council Recap – incredible meeting, main focus: plans for West renovation. The Waldo Road storage space now officially belongs to the Library. Handout of Library Phasing Options was distributed, showing four possible plans involving book/staff relocations. Possible locations for 1B Transplant phase: the old K-Mart on 13th Street or the basement of the Oaks Mall. If staff and books relocate to an “alternate library” facility, there would have to be public space, bathrooms, safety considerations, etc. We have signed-off not to use too much brick (not enough money). The compromise is that we will have lots of windows. We have 5 million volumes to store. If money runs out, then everything is on hold – but we will be breaking ground come November. We’re getting closer to the deadline, but have no plan to implement. There is a Cox and GRU POP close to the Oaks Mall. Jan Swanbeck has insured that Documents will move out of West to MSL. We had to rush the estimates for the wiring. Some of the money for construction in West will pay for the conduit in MSL. Documents will have to make the physical move to MSL. The walk-up section on the second floor of MSL will need to be even bigger. At first it was going to have 60 workstations, it has now been cut to 36. They also want wireless access. None of this can happen until the metal file cabinets are moved. Compact shelving must also move. There is still dispute regarding the walls. We want to do all this in the fall. The book movers do not have enough people in the area – those that commute in will need a place to stay – there would not much available on football weekends, so this has to be scheduled for a time when there are no home games 2-3 weeks in a row. Looks like this may happen during the first week of October. Steve Shorb will present the next public overview of the plans on Wednesday. LibQual Survey – patrons asked to rate what Library services are vs. what they should be. Have had fewer respondents this year – less than 300 – statistically irrelevant.
Review Minutes of 07/21/03
  1. Copier part is not in yet.
  2. FreezeOut has not been downloaded yet.
Ongoing Systems Updates
Equipment Received/Installed/Repaired
  1. Regular ordering – purchase cards are working.
  2. Major water pipe break in West. Hallway and 148 classroom were flooded, as well as some of the HSSS staff area.
  3. Lightning hit a switch over in AFA late Friday afternoon. Damage is where the fiber comes into the switch. There are no ports on that switch. A PC probably took a hit and it burned through. Bobbie replaced the switch. Need to order another switch for the shelf.
  1. VOIP – for Documents in MSL basement. The phones are network switches – use POE – phone to network jack – requires power from the jack. If the machines are down, the phones will be, too. It will require special switches – ours can’t do that. Net services will be handling VOIP.
  2. New server status - Docutek – Will finished, tested – it works fine. He turned it over to Docutek, after that, the modem did not work. To install a new modem, Will had to uninstall, remove the old modem, then reinstall the new one – basically start again from scratch. It is now back up and running. Docutek wants to do all of the updates and installs remotely. It has VNC (virtual network control) – you can actually watch on the console - Bill observed them changing the password. Ironically, they called because they could not get in – they’d had forgotten what they changed the password to and Bill had to reset it.
Software Projects Status
  1. Problem with the Verify Patron Privacy Notice – people were able to type over it. It is now Read Only. Its location can be moved, but it returns to the original position after each logoff. The new version (3.3) is housed in the AFLUSER folder.
  2. Cynthia and Winston have been working on SysRequest to upload documents. Need to figure out where to put them and make them available.
  3. Hilton is working on FilmLog (Preservation). They have issues with the interface.
  4. Book Film Replacement – ongoing.
  5. GIS – Joe is back – he is pleased with what Ying has done - had Joe convince Erich he was pleased – good support.
  6. Martha – question on OPS Work Study report (last year’s totals) for Personnel.
Web Projects
  1. New candidate prospects – one in the pool.
Student and Department Happenings
  1. Bob Leemer will be here tomorrow to work on the cartridge cleaner. It has a problem at the take-up spool.
  2. Students re-verified that all of the tape library volsers are in the correct rack/slot space following the recent moves.

Back to the Systems Homepage

Last updated August 4, 2003
by Debra Fetzer